OpenStudy (anonymous):

There is a number that is 8 less than the sum of 20 + 12. Find the number. Let n = the number.

OpenStudy (anonymous):

Which equation should be used to solve this problem?

OpenStudy (anonymous):

20+12=32 32-8=24

OpenStudy (saifoo.khan):

Working..?

OpenStudy (anonymous):

n-8=(20-12) need something like that

OpenStudy (anonymous):

n=(20+12)-8 was the right answer

OpenStudy (anonymous):

@aussy123 was the closest thanks:)

OpenStudy (anonymous):

no prob